OpenStudy (anonymous):

Help how do I evaluate f(a+h)-f(a)/h for f(x)=x^2 +3???

OpenStudy (anonymous):

@amistre64 @Paynesdad

OpenStudy (amistre64):

plug in f(a) and f(a+h)

OpenStudy (anonymous):

Just the difference quotient, or are you taking a limit to find the derivative? In any case, given that \(f(x)=x^2+3\), you have \(f(a)=a^2+3\) and \(f(a+h)=(a+h)^2+3\). So, the difference quotient is \[\frac{f(a+h)-f(a)}{h}=\frac{((a+h)^2+3)-(a^2+3)}{h}\] Then simplify.

OpenStudy (anonymous):

so you have to plug in x^2+3 every time there is an f? @SithsAndGiggles

OpenStudy (anonymous):

You have to plug in whatever is in the parentheses: \[f(\cdots)=(\cdots)^2+3\] In this case, \(\cdots\) represents \(a\) and \(a+h\).

OpenStudy (anonymous):

ok got it thanks!